25ml of a solution containing HCl and H2SO4 required 10ml of a 1N NaOH solution for  neutralization. 20ml of same acid mixture on being treated with an excess of AgNO3 gives 0.1435g of AgCl. The normality of the HCl and normality of the H2SO4 are 

Grade:12th pass

3 Answers

Arun
5 years ago
 
Dear Pawan
 
I am getting my aswer as 
0.05N & 0.35 N
 
Let me know if thisi is correct then I will share my solution.
 
Regards
Arun (askIITians forum expert)
Pawan joshi
5 years ago
Yes, the answer is correct sir
Devansh
4 years ago
Let normality be N1 and N2
Meq HCl +Meq H2SO4 =Meq NaOH
25*N1 +25*N2 =10*1
=> N1+N2 =0.4
By POAC
moles of AgCl =moles of Cl 
(20*N1)/1000 =0.1435/143.5 =10^{-3}
N1=0.05N
N2=0.35N
moles of AgCl =